# MalMixer: Few-Shot Malware Classification with Retrieval-Augmented Semi-Supervised Learning

### v2

@@ VII-A RQ1: Saturation Analysis
In this subsection, we measure how well MalMixer retains adequate performance when its augmentation pipelines is provided with varying number of labeled data for reference. We use a standardized semi-supervised set-up over the BODMAS-20 dataset, and report accuracy, precision, recall, and F1 (all macro) as key criteria.
@@
Result for Saturation Analysis Table IV shows that substantially decreasing the volume of labeled samples does not dramatically undermine the performance of MalMixer. We found a labeled set consisting of $1\%$ of all data balances the number of labeled samples with overall model performance. Furthermore, MalMixer seems unbiased across a variety of families from different paradigms, as shown in Table IX. We note that MalMixer’s pervasive use of regularization likely means the model is underfit and thus cannot reach the performance of a fully-supervised model (e.g., ResNet-12), which is the gold standard for classification and serves as a reference. Nonetheless, our approach can provide adequate classification accuracy with very few labeled samples.
VII-B RQ2: Model Comparison
@@
We further evaluate MalMixer in few-shot settings on BODMAS-20 and MOTIF-50 by comparing against six other baseline models trained on static features: Pseudo-Label, VAT, VAT (aug), MixUp, MixMatch, and MixMatch (Gaussian). We use accuracy, precision, recall, and F1 (all macro) as criteria.
@@
Result for Model Comparison Table V compares MalMixer against baseline models in few-shot settings with $0.1\%$ labeled data on BODMAS-20 and $1\%$ labels on MOTIF-50, where most families are identified by one-shot labels.
@@
On BODMAS-20, MalMixer achieves a relative accuracy improvement of at least $5.60\%$ compared to all but one of the baseline models. The only model not outperformed by such a margin is VAT (aug), which employs our malware augmentation method. In fact, VAT (aug), on either BODMAS-20 or MOTIF-50, sees a relative improvement of $6.23\%$ or $7.49\%$ over VAT. This demonstrates our augmentation technique also effectively enhances other compatible semi-supervised frameworks. Moreover, MalMixer surpasses MixMatch with Gaussian noise augmentation by a relative margin of $7.51\%$ on BODMAS-20 and $15.94\%$ on MOTIF-50, setting our proposed augmentation pipeline apart from semantics-agnostic, statistical augmentation approaches.
@@
Figure 8 illustrates that the accuracy of MalMixer and the baseline models converge as the number of labeled data increases. We note that all models perform significantly worse on MOTIF-50 due to the intrinsic complexities and small sizes of malware classes in MOTIF, where fully supervised state-of-the-art models only achieved $48.7\%$ and $72.4\%$ accuracy [26].
@@
Figure 8: Comparison of MalMixer to baseline methods in few-shot malware classification settings on BODMAS-20. MalMixer outperforms all models with fewer (i.e. towards the leftmost end of the x-axis) labeled data, but performance converges with a higher proportion of labeled data.
TABLE VI: Comparison of MalMixer against other semi-supervised methods on malware classification of BODMAS-Temporal.
@@
The few-shot performance of MalMixer in results for RQ1–2 showcases the effectiveness of our augmentation approach in overcoming small labeled datasets.
VII-C RQ3: Temporal Analysis
@@
In this setting, all malware families of interest are represented in train set, but the evaluation samples follow a different distribution of families. We evaluate our model against the BODMAS-Temporal dataset, and report accuracy, precision (macro), recall (macro), and F1 (macro) as key criteria.
@@
MalMixer effectively improves classification of new samples from previously identified families. Our augmentation helps to generalize classifiers in the face of shifting distribution of in-the-wild malware over time.
VII-D RQ4: Leave-Out Analysis
Next, we consider a scenario in which a new, previously unlabeled malware families must be classified. In practice, new malware may appear over time as attackers find new zero-day exploits. In the leave-out setting, some families are originally unidentified in training. Using the BODMAS-LeaveOut dataset, we evaluate the performance improvement of MalMixer when one labeled sample from each previously ‘unidentified’ family is introduced to training. We employ accuracy, precision, recall, and F1 (all macro) as criteria.
Result for Leave-Out Analysis Table III illustrates improvements in MalMixer’s performance when one-shot instances from formerly ‘unidentified’ families become labeled. In the set-up with 21 malware families, the introduction of merely one additional label for the originally ‘unidentified’ malware family (constituting about $5\%$ of all data) increases MalMixer’s accuracy by $4.57\%$ through straightforward retraining, which can be completed within an hour. In more challenging scenarios, where 10 families (accounting for roughly $44\%$ of all data) in the test set are deemed ‘unidentified’ in training, the addition of just 10 labels can increase the model’s accuracy by a substantial $21.53\%$.
Leave-out results imply MalMixer’s performance improves effectively through a quick retraining with minimal labeling of previously unidentified families.
VII-E Ablation Studies
Last, we conduct ablation studies to understand how each component of our augmentation pipeline contributes to MalMixer’s performance. We evaluate this using the BODMAS-Temporal set-up, maintaining the training set while using all 2020 data for testing. Table VIII displays the performance decline under three conditions: (1) without the alignment-based malware selection step, where we retrieve any real set of non-interpolatable features representative of the mixed instance, disregarding its compatibility with the mutated interpolatable features; (2) without both the retrieval-based malware matching and alignment-based malware selection steps, where we eliminate the use of embeddings altogether and perform augmentation by directly interpolating two malware instances’ original features without semantics awareness; and (3) without our mixing-based augmentation, where we solely rely on Gaussian noise augmentation.
@@ VIII Discussion
Interpretation and Implication of Findings The collective findings from RQ1–4 suggest that MalMixer can effectively improve few-shot malware family classification performance given an influx of new malware samples by being able to (1) accurately classify malware families with limited labels; (2) generalize to evolved instances from previously identified families without substantial new labeling; and (3) effectively retrain to detect new families with minimal new labels. These findings imply that, in contrast to the semantics-agnostic imaged-based augmentation, malware augmentation pipelines can benefit from domain-knowledge-awareness achieved through augmenting different sets of malware features differently, as well as retrieval and alignment techniques capturing feature semantics. Furthermore, equipped with such domain-knowledge-aware augmentation approaches, semi-supervised frameworks analogous to those implemented in MalMixer can be successfully extended to malware family classification settings. Intuitively, MalMixer works because few-shot samples sparse within the true distribution of malware lead to difficulty drawing decision boundaries. Our approach uses domain knowledge to more faithfully mimic true malware family feature distribution, narrowing down decision boundaries. In practice, we speculate that MalMixer can help alleviate the extensive manual labor required to classify a large dataset.
@@
Limitations and Future Work: We note that while our augmentation approach is domain-knowledge-aware, it does not guarantee that augmented pseudo-samples can be reverse-engineered into an exact real-world malware executable. That said, augmentation methods that fully preserve the semantics of a malware executable often require semantics-aware transformations on malware binaries [20], or possibly another diffusion step to recover a real binary out of our augmented feature representations, both of which can be time and resource-intensive. In our approach, we made a deliberate trade-off between some semantic validity and operational efficiency. We contain our claim of “semantic-awareness” to the use of domain knowledge in order to become aware of what feature interpolations more/less likely change the underlying malware family semantics. This is an improvement over conventional image-based augmentations (which flip random bits without understanding any consequences), and has been shown effective with improved downstream ML classifiers’ performance (Section VII) and the fact that augmented feature representations fall back into the clusters of ground-truth families (Section IV). From a downstream ML model’s perspective, our augmentation approach, although not perfectly semantics-preserving, already better mimics the underlying distribution of each family in the feature space.
@@
Moreover, MalMixer currently accepts malware instances represented by features obtained through static analysis. While static-feature-based classification is a crucial research domain [43] and is often more effective than end-to-end classification that consumes entire malware binaries [42], future work can investigate efficient and semantics-aware augmentation methods that target dynamic analysis features. Nonetheless, the benefits of static analysis features, such as efficiency, scalability, and safety, hold substantial value for numerous practical applications, and are readily extracted from binaries. Additionally, static features may be susceptible to code obfuscation. While we verify that MalMixer already outperforms baseline models given few-shot samples of newly obfuscated samples, as shown in Section E in the Appendix, future research can still improve static feature-based classifiers’ robustness against obfuscation.
Lastly, our model relies on a closed set assumption, whereby it can only classify malware samples into families identified during training. While this is a shared limitation among traditional ML algorithms, MalMixer already aims to mitigate this issue through few-shot learning: with various existing techniques able to identify the emergence of new families [58], MalMixer effectively works as a follow-up procedure after new families have been identified. When novel malware families appear, MalMixer can be quickly retrained to classify larger datasets after only a few samples become labeled, thereby reducing the number of labeled samples needed to understand an emerging family.
IX Conclusion
@@
We present MalMixer, a few-shot malware family classifier unifying a novel, lightweight, domain-knowledge-aware malware augmentation approach and a semi-supervised framework previously unexplored in malware classification contexts. Our experiments show that our classifier outperforms others in few-shot malware classification settings. This research verifies the feasibility and effectiveness of lightweight, domain-knowledge-aware malware feature augmentation methods, which prove valuable when few labeled samples are available. We further showcase the capabilities of the implemented semi-supervised framework to bolster the effect of domain-knowledge-aware malware augmentation and thereby improve few-shot malware family classification performance.
Appendix A Formulation of the Label Guessing, Average & Sharpening, and MixUp Steps of MalMixer
In MalMixer’s semi-supervised framework, training data first undergoes data augmentation to generate synthetic samples (Section III). Using the same notation as before, we denote the original labeled data $\mathcal{X}$, their labels $\mathcal{Y}$, and the original unlabeled data $\mathcal{U}$. We denote the data augmentation process $\mathcal{A}$ and $\mathcal{X}^{\prime}=\mathcal{X}\cup\mathcal{A}(\mathcal{X}),\mathcal{U}^{\prime}=\mathcal{U}\cup\mathcal{A}(\mathcal{U})$.
Label Guessing, Average & Sharpening Upon obtainment of $\mathcal{U}^{\prime}$, we guess the pseudo labels for $\mathcal{U}^{\prime}$, denoted $\mathcal{Y}_{\mathcal{U}}$:
$\mathcal{Y_{\mathcal{U}}}=\frac{P(Y|\mathcal{U})+P(Y|\mathcal{A}(\mathcal{U}))}{2},$
where $P$ is the base classifier (e.g. ResNet on a batch) and $\mathcal{Y_{\mathcal{U}}}$ is the pseudo-label generated through averaging the classifier’s predictions between $\mathcal{U}$ and $\mathcal{A}(\mathcal{U})$. Sharpening is then applied to the average prediction. For simplification, in the following formula, we let $\mathcal{X^{\prime}}=(\mathcal{X^{\prime}},\mathcal{Y}\cup\mathcal{Y})$, and $\mathcal{U^{\prime}}=(\mathcal{U^{\prime}},\mathcal{Y_{\mathcal{U}}}\cup\mathcal{Y_{\mathcal{U}}})$, meaning $\mathcal{X^{\prime}}$ and $\mathcal{U^{\prime}}$ become tuples containing both data and label.
MixUp Upon obtaining the pseudo labels for the augmented unlabeled data, we obtain a fully labeled dataset, denoted as $\mathcal{W}=\text{shuffle}({\mathcal{X}^{\prime}\cup\mathcal{U}^{\prime}})$. Subsequently, MixUp [37] is applied to $\mathcal{X}^{\prime}$, $\mathcal{U}^{\prime}$, and $\mathcal{W}$ to obtain $\mathcal{X^{\prime}}=M(\mathcal{X^{\prime}},\mathcal{W_{X^{\prime}}}),\mathcal{U^{\prime}}=M(\mathcal{U^{\prime}},\mathcal{W_{U^{\prime}}})$. Here, $\mathcal{W_{X^{\prime}}},\mathcal{W_{U^{\prime}}}$ are disjoint subsets of $\mathcal{W}$ with $|\mathcal{W_{X^{\prime}}}|=|\mathcal{X}^{\prime}|$ and $|\mathcal{W_{U^{\prime}}}|=|\mathcal{U}^{\prime}|$; $M$ represents the linear interpolation of both features and labels (one-hot).
To train a base classifier, the loss for malware prediction is
$\mathcal{L_{X^{\prime}}}$ $=\frac{1}{|\mathcal{X^{\prime}}|}\sum_{x,y\in\mathcal{X^{\prime}}}{CE(y,p(y|x))},$
where $p$ denotes the base classifier (for a single datapoint), $CE$ represents the cross-entropy loss function. The final loss is
$\mathcal{L}=\mathcal{L_{X^{\prime}}}+\lambda\mathcal{L_{U^{\prime}}},~{}\lambda\text{ is a hyperparamter}.$ (5)
Appendix B Implementation Details

### v4

@@ 7.1 RQ1: Saturation Analysis
In this subsection, we measure how well MalMixer retains adequate performance when its data augmentation pipelines is provided with varying number of labeled data for reference. We use a standardized semi-supervised set-up over the BODMAS-20 dataset, and report accuracy, precision, recall, and F1 (all macro) as key criteria.
@@
Result for Saturation Analysis Table IV shows that substantially decreasing the volume of labeled samples does not dramatically undermine the performance of MalMixer on BODMAS-20. We found a labeled set consisting of $1\%$ of all data balances the number of labeled samples with overall model performance on this dataset. Furthermore, MalMixer seems unbiased across a variety of families from different paradigms, as shown in Table XI in the Appendix. We note that MalMixer’s pervasive use of regularization likely means the model is underfit and thus cannot reach the performance of a fully-supervised model (e.g., Fully Connected ResNet), which is the gold standard for classification and serves as a reference. Nonetheless, our approach can provide adequate classification accuracy with very few labeled samples.
Furthermore, the supervised performance of our chosen base classifier, FC-ResNet (Appendix A), as shown in Table IV, demonstrates that FC-ResNet possesses sufficient expressiveness to serve as the base classifier within our semi-supervised framework. Its performance remains competitive with LightGBM [59], a state-of-the-art model used in the original BODMAS experiments [24]. See more discussion on FC-ResNet’s expresiveness in Appendix C.
7.2 RQ2: Model Comparison
@@
We further evaluate MalMixer in near one-shot settings on BODMAS-20 and MOTIF-50 by comparing against seven other baseline models trained on static features: Pseudo-Label, VAT, VAT (aug), MixUp, MixMatch, MixMatch (Gaussian), and MORSE. We use accuracy, precision, recall, and F1 (all macro) as criteria.
@@
Result for Model Comparison Table V compares MalMixer against baseline models with $0.1\%$ labeled data on BODMAS-20 and $1\%$ labels on MOTIF-50, where most families are identified by one-shot labels.
@@
On BODMAS-20, MalMixer achieves a relative accuracy improvement of at least $5.41\%$ compared to all but one of the baseline models. The only model not outperformed by such a margin is VAT (aug), which employs our malware data augmentation method. In fact, VAT (aug), on either BODMAS-20 or MOTIF-50, sees a relative improvement of $6.23\%$ or $7.49\%$ over VAT. This demonstrates our augmentation technique also effectively enhances other compatible semi-supervised frameworks. Moreover, MalMixer surpasses MixMatch with Gaussian noise data augmentation by a relative margin of $7.51\%$ on BODMAS-20 and $15.94\%$ on MOTIF-50, setting our proposed augmentation pipeline apart from semantics-agnostic, statistical augmentation approaches.
MalMixer outperforms MORSE, likely due to fundamental differences in design objectives. MORSE is tailored for scenarios with more than $10\%$ labeled data, making many of its elegant mechanisms impractical in few-shot settings. Furthermore, MORSE’s proposed data augmentation, although malware-specific, still largely disregards the underlying semantics behind features. We provide further discussion on these points in Appendix I.
@@
Figure 8 illustrates that the accuracy of MalMixer and the baseline models converge as the number of labeled data increases. We note that all models perform significantly worse on MOTIF-50 due to the intrinsic complexities, noisy labels [47], and small sizes of malware classes in MOTIF.
@@
Figure 8: Comparison of MalMixer to baseline methods in few-shot settings on BODMAS-20. MalMixer outperforms all models with fewer (i.e. towards the leftmost end of the x-axis) labeled data, but performance converges with a higher proportion of labeled data.
TABLE VI: Comparison of MalMixer against baseline methods on malware classification of BODMAS-Temporal.
@@
The few-shot performance of MalMixer in results for RQ1–2 showcases our augmentation approach’s effectiveness in overcoming small labeled datasets.
7.3 RQ3: Temporal Analysis
@@
In this setting, all malware families of interest are represented in train set, but the evaluation samples follow a different distribution of families. We evaluate our model against the BODMAS-Temporal dataset, and report accuracy, precision, recall, and F1 (all macro) as key criteria.
@@
MalMixer effectively improves classification of new samples from previously identified families. Our proposed data augmentation helps to generalize classifiers in the face of shifting distribution of in-the-wild malware over time.
7.4 RQ4: Leave-Out Analysis
Next, we consider a scenario in which new, previously unlabeled malware families must be classified. In practice, new malware may appear over time as attackers find new zero-day exploits. Classifying these unseen-family instances using MalMixer requires retraining the model after experts manually provide labels for each new family. Using the BODMAS-LeaveOut dataset, we evaluate the performance improvement of MalMixer when as few as one labeled sample from each previously ‘unidentified’ family is introduced to training. We employ accuracy, precision, recall, and F1 (all macro) as criteria.
Result for Leave-Out Analysis Table VII illustrates improvements in MalMixer’s performance when one-shot instances from formerly ‘unidentified’ families become labeled. In the set-up with 21 malware families, the introduction of merely one additional label for the originally ‘unidentified’ malware family (constituting about $5\%$ of all data) increases MalMixer’s accuracy by $4.57\%$ through straightforward retraining, which can be completed within an hour. In more challenging scenarios, where 10 families (accounting for roughly $44\%$ of all data) in the test set are deemed ‘unidentified’ in training, the addition of just 10 labels can increase the model’s accuracy by a substantial $21.53\%$.
Leave-out results imply MalMixer’s performance improves effectively through a quick retraining with minimal labeling of previously unidentified families, reducing the amount of manual labeling required to classify a new family.
7.5 Ablation Studies
Last, we conduct ablation studies to understand how each component of our data augmentation pipeline contributes to MalMixer’s performance. We evaluate this using the BODMAS-Temporal set-up, maintaining the training set while using all 2020 data for testing. Table VIII displays the performance decline under three conditions: (1) without the alignment-based malware selection step, where we retrieve any real set of non-interpolatable features representative of the mixed instance, disregarding its compatibility with the mutated interpolatable features; (2) without both the retrieval-based malware matching and alignment-based malware selection steps, where we eliminate the use of embeddings altogether and perform data augmentation by directly interpolating two malware instances’ original features without semantics awareness; and (3) without our mixing-based data augmentation, where we solely rely on Gaussian noise augmentation.
@@ 8 Discussion
Interpretation and Implication of Findings The collective findings from RQ1–4 suggest that MalMixer can effectively improve few-shot malware family classification performance given an influx of new malware samples by being able to (1) accurately classify malware families with limited labels; (2) generalize to evolved instances from previously identified families without substantial new labeling; and (3) effectively retrain to detect new families with minimal new labels. These findings imply that, in contrast to the semantics-agnostic image-based data augmentation, malware data augmentation pipelines can benefit from domain-knowledge-awareness achieved through augmenting different sets of malware features differently, as well as retrieval and alignment techniques capturing feature semantics. Furthermore, equipped with such domain-knowledge-aware data augmentation approaches, semi-supervised frameworks analogous to those implemented in MalMixer can be successfully extended to malware family classification settings. Intuitively, MalMixer works because few-shot samples sparse within the true distribution of malware lead to difficulty drawing decision boundaries. Our approach uses domain knowledge to more faithfully mimic true malware family feature distribution, narrowing down decision boundaries. In practice, we speculate that MalMixer can help alleviate the extensive manual labor required to classify a large dataset.
@@
Limitations and Future Work: We note that while our data augmentation approach is domain-knowledge-aware, it does not guarantee that augmented pseudo-samples can be reverse-engineered into an exact real-world malware executable. That said, data augmentation methods that fully preserve the semantics of a malware executable often require semantics-aware transformations on malware binaries [20], or possibly another diffusion step to recover a real binary out of our augmented feature representations, both of which can be time and resource-intensive. In our approach, we made a deliberate trade-off between some semantic validity and operational efficiency. We contain our claim of “semantic-awareness” to the use of domain knowledge in order to become aware of what feature interpolations more/less likely change the underlying malware family semantics. This is an improvement over conventional image-based data augmentations (which flip random bits without understanding any consequences), and has been shown effective with improved downstream ML classifiers’ performance (Section 7) and the fact that augmented feature representations fall back into the clusters of ground-truth families (Section 4). From a downstream ML model’s perspective, our data augmentation approach, although not perfectly semantics-preserving, already better mimics the underlying distribution of each family in the feature space.
Currently, MalMixer is evaluated only on BODMAS and MOTIF, both within the Windows PE malware domain. Future work could extend its applicability to other domains (e.g. Android malware), enhancing generalizability. In principle, adapting our approach to another domain involves manually re-defining the operations in Appendix G (which took approximately five hours to complete by one author familiar with malware classification).
@@
Moreover, MalMixer currently accepts malware instances represented by features obtained through static analysis. While static-feature-based classification is a crucial research domain [43] and is often more effective than end-to-end classification that consumes entire malware binaries [42], future work can investigate efficient and semantics-aware augmentation methods that target dynamic analysis features. Nonetheless, the benefits of static analysis features, such as efficiency, scalability, and safety, hold substantial value for numerous practical applications, and are readily extracted from binaries.
Additionally, static features are susceptible to code obfuscation. To the end, Appendix F also evaluates MalMixer on packed malware samples. While this experiment is limited by noisy family labels, it demonstrates negative impacts of packing on MalMixer’s performance, likely due to the weakening of our $k$-NN and clustering assumptions in Sections 1 and 3.3. Nevertheless, MalMixer remains among the top-performing methods, and our semantics-aware data augmentation pipeline is at least competitive with Gaussian-noise-based, image-inspired data augmentation techniques under obfuscated conditions. These findings highlight the need for future research to enhance the robustness of static feature-based classifiers against obfuscation.
We also limited our saturation analysis to BODMAS-20. Appendix B conducts an additional saturation analysis on MOTIF-50. Results show that MalMixer achieves satisfactory performance in few-shot settings on MOTIF-50, although such satisfactory performance requires a higher percentage of labeled data compared to BODMAS-20. We hypothesize that the high label noise in MOTIF-50 renders it less suitable for an accurate saturation analysis, and thus omitted it from the main text. Future work aimed at benchmarking models’ performance degradation with decreasing numbers of labels via saturation analysis could benefit from additional Windows PE datasets featuring more carefully curated family labels.
In addition, we only evaluated MalMixer on downsampled versions of BODMAS and MOTIF, focusing primarily on the top-N malware families. MalMixer is designed to expedite classification for large families. In real-world settings, the influx of malware data is typically characterized by a long-tail distribution with many small families. As experts generally have the capacity to promptly label only the larger, more prevalent families, the labeled pool naturally skews toward these major families. Our evaluation aligns the unlabeled data distribution and test distribution with that of the labeled samples, thus only evaluating MalMixer’s performance on on large families. Appendix E further evaluates MalMixer’s performance on non-downsized datasets, in presence of more and smaller families. Results show that MalMixer’s performance on large families remains strong, even with small unlabeled families falsely classified into these larger families. When only few-shot labels for large families are available, preliminarily misclassifying unlabeled small families into larger families can still provide insightful information about the characteristics of the unlabeled samples and allow engineers to prioritize further analysis. Appendix E also shows MalMixer’s performance naturally degrades with more families — a well-known challenge in multi-way classification, as also observed with the supervised baseline. We acknowledge this limitation of our work, and leave enhancing robustness against an increasing number of classes as important future research.
Practical Implications Our model operates on a closed-wold assumption, whereby it can only classify malware samples into families identified during training. This assumption aligns with real-world scenarios where we wish to train a classifier from scratch given an influx of unlabeled dataset. Once experts provide few-shot labels, likely for large families, MalMixer’s few-shot capacity allows for prompt classifications of a largely unlabeled dataset, as discussed above.
In some other cases, we may wish to apply a pre-trained MalMixer to a newly arrived test set containing malware families not seen during training, resulting in an open-world setting. MalMixer is not designed or trained for a reject option for potentially unseen families. Instead, existing techniques [57] (e.g. fully-supervised models explicitly designed and trained with decision threshold) can detect unknown families, almost surely outperforming few-shot MalMixer on the rejection task. Once existing techniques reject new families and experts manually identify one labeled instance per new family, the problem space becomes closed-world, and MalMixer becomes practically useful for what happens next. This scenario is simulated in Section 7.4, where MalMixer’s few-shot capability allows it to quickly retrain to accurately classify all instances of these new families using few labeled samples, thereby reducing the number of labeled samples needed to understand an emerging family.
9 Conclusion
@@
We present MalMixer, a few-shot malware family classifier unifying a novel, lightweight, domain-knowledge-aware malware data augmentation approach and a semi-supervised framework previously unexplored in malware contexts. Our experiments show that our classifier outperforms others in few-shot malware classification settings. This research verifies the feasibility and effectiveness of lightweight, domain-knowledge-aware data augmentation methods for malware features, which prove valuable when few labeled samples are available. We further showcase the capabilities of the implemented semi-supervised framework to bolster the effect of domain-knowledge-aware malware data augmentation and thereby improve few-shot malware family classification performance.
Acknowledgment
We acknowledge partial support from the NSA under grant H98230-23-C-0279, the NSF under grant DGE2312057, and an Amazon Research Award. The views expressed here do not necessarily reflect those of the US Government.
Appendix A Implementation Details